OpenStudy (spring98):

The table and the graph below each show a different relationship between the same two variables, x and y: How much more would the value of y be on the graph than its value in the table when x = 12? 20 30 60 70

OpenStudy (spring98):

@iGreen

OpenStudy (spring98):

i think its 30 but im not sure!

OpenStudy (spring98):

@iGreen

OpenStudy (igreen):

Table increases by 25..the graph increases by 30. 25 * 12 = ? 30 * 12 = ?

OpenStudy (spring98):

ohh 300 and 360

OpenStudy (spring98):

@iGreen

OpenStudy (igreen):

Yep, now subtract. 360 - 300 = ?

OpenStudy (spring98):

60

OpenStudy (spring98):

so it would be 60 right?

OpenStudy (igreen):

Yep

OpenStudy (spring98):

k thanks